### v1

# Simultaneous Khintchine theorem on manifolds in positive characteristics: convergence case

Abstract.
In this article, we prove the convergence case of Khintchine’s theorem for analytic nonplanar manifolds over local fields of positive characteristic, in the setting of simultaneous Diophantine approximation. Our approach is based on the method of counting rational points near manifolds developed by Beresnevich and Yang [BY23]. The results obtained here extend the work of Beresnevich and Yang [BY23], and Beresnevich and Datta [BD25], to the function field setting. In the course of the proof, we also establish several new results in the geometry of numbers over function fields, which are of independent interest.
Key words and phrases:
Diophantine approximation, Khintchine Theorem
2020 Mathematics Subject Classification:
Primary 11J13, 11J61, 11J83; Secondary 11J70,37A17,22E40
1. Introduction
Diophantine approximation is a central topic in number theory, which deals with the effective density of rational numbers within real numbers and its higher-dimensional analogue. The Khintchine-Groshev theorem is a foundational result in metric Diophantine approximation. Let $\psi:(0,\infty)\rightarrow(0,1)$ be a approximating function, i.e., a non-increasing function such that $\psi(x)\rightarrow 0$ as $x\rightarrow\infty.$ For a fixed $\bm{\theta}\in\mathbb{R}^{n}$ with $n\in\mathbb{N}$, the following set is an object of interest for a long time
$\mathcal{S}_{n}^{\bm{\theta}}(\psi):=\left\{\mathbf{y}\in\mathbb{R}^{n}:\left\|\mathbf{y}-\frac{\mathbf{p}+\bm{\theta}}{q}\right\|<\frac{\psi(q)}{q}\,\,\text{for i.m. }(\mathbf{p},q)\in\mathbb{Z}^{n}\times\mathbb{N}\right\},$
where $\|\cdot\|$ denotes the sup norm and ‘i.m’ reads as ‘infinitely many’. The points $\mathbf{y}$ lying in $\mathcal{S}_{n}^{\bm{\theta}}(\psi)$ are usually referred to as $(\psi,\bm{\theta})$-approximable. The case $\bm{\theta}=\mathbf{0}$ is called homogeneous setting, and the points of $\mathcal{S}_{n}^{\mathbf{0}}(\psi)$ are called $\psi$-approximable. The form of approximation described above concerns how well one can approximate points of $\mathbb{R}^{n}$ by rational points, and it is commonly referred to as the simultaneous form of approximation. In contrast, there is another widely studied type of approximation, where one asks how close a point in $\mathbb{R}^{n}$ is to a rational hyperplane; this is known as the dual form of approximation. To begin with, we recall the inhomogeneous version of classical Khintchine’s theorem (see [Khi26, Gro38] and [AR23, §1.1]).
Theorem 1.1 (The Inhomogeneous Khintchine theorem).
Given any approximating function $\psi$ and $\bm{\theta}\in\mathbb{R}^{n}$
$\mathcal{S}_{n}^{\bm{\theta}}(\psi)=\begin{cases}\text{Lebesgue null}&\text{if }\displaystyle\sum_{q=1}^{\infty}\psi(q)^{n}<\infty\\ \text{Lebesgue full}&\text{if }\displaystyle\sum_{q=1}^{\infty}\psi(q)^{n}=\infty.\end{cases}$ (1.1)
The situation becomes more delicate when examining the extent to which embedded submanifolds of $\mathbb{R}^{n}$ inherit the Diophantine properties prevalent in $\mathbb{R}^{n}$. Establishing an analogue of Khintchine’s theorem for manifolds is closely linked to the challenging problem of counting rational points near the manifold (for example, see [Ber02, Ber12, BVVZ21, SST23, BD25]). For analytic non-degenerate manifolds of dimension $d\geq 2$, the divergence case was established by Beresnevich [Ber12], and later extended to non-analytic curves in [BVVZ21], both employing the ubiquity framework alongside Kleinbock–Margulis’s quantitative non-divergence estimates [KM98]. For non-degenerate manifolds in $\mathbb{R}^{n}$, the convergence case was resolved by Beresnevich and Yang [BY23] (see also [SST23]), and Khintchine’s theorem in full generality for simultaneous approximation was recently settled by Beresnevich and Datta [BD25]. More recently, Datta [Dat24] extended the result in Beresnevich and Yang by proving a quantitative simultaneous Khintchine theorem over manifolds in $\mathbb{R}^{n}$.
Unlike the simultaneous form, the dual form of Diophantine approximation has long been well understood, beginning with Kleinbock and Margulis’s resolution of the Baker-Sprindžuk conjecture [KM98]. A complete dual Khintchine theorem for non-degenerate manifolds was later established in both homogeneous [BKM01, BBKM02] and inhomogeneous settings [BBV13].
In recent years, there has been growing interest in Diophantine approximation over local fields of positive characteristic. Diophantine approximation over function fields concerns the quantitative study of approximating Laurent series by rational functions and their higher-dimensional analogues. In this context, Mahler developed the geometry of numbers in the seminal paper [Mah41], which provides a straightforward route to proving the analogue of Dirichlet’s theorem. The Khintchine theorem in positive characteristic was proved by de Mathan [dM70], later extended by Kristensen to systems of linear forms and to the analysis of Hausdorff dimensions of exceptional sets [Kri03]. Further advances include a multiplicative Khintchine–Groshev theorem [AGP12] and extensions to imaginary quadratic function fields [GR15]. More recently, Chao Ma and Wei-Yi Su established inhomogeneous versions of the Khintchine and Jarník–Besicovitch theorems [MS08], and Kristensen derived asymptotic formulas for inhomogeneous linear systems, yielding inhomogeneous forms of the Khintchine–Groshev and Jarník theorems [Kri11]. For further developments in the inhomogeneous theory, see [KN11, Fuc10].
In the metric theory over manifolds in this setting, the analogues of the Baker-Sprindžuk conjecture had been proved by Ghosh [Gho07]. Subsequently, Ganguly and Ghosh proved the inhomogeneous Sprindžuk conjecture in [GG20]. Recently, Das and Ganguly proved a complete inhomogeneous Khintchine-Groshev type theorem for nonplanar manifolds over function fields in the dual setting [DG22]. For additional results and recent developments on Diophantine approximation in function fields, see [GG19, Gan22, DG24, Ban25, AK25, DX24, KLP23].
In this paper, we prove the convergence case of Khintchine’s theorem for analytic non-planar manifolds over local fields of positive characteristic, extending the earlier works of Beresnevich and Yang [BY23], and Beresnevich and Datta [BD25] to the setting of function fields. Although our approach is based on the technique of Beresnevich and Yang [BY23] for counting rational points near manifolds, the positive characteristic setup poses many challenges, which we have explained in §1.4. Apart from this, we have also proved some new results in the geometry of numbers over function fields (see §3 and §6 ), which are of independent interest. We begin by introducing the function field framework.
1.1. The Function Field Setting
We begin with the field of rational functions $\frac{P}{Q}\in\mathbb{F}_{q}(T)$, where $q=p^{b}$ for some prime $p$ and $b\in\mathbb{N}.$ We define a non-archimedean absolute value on $\mathbb{F}_{q}(T)$ as follows. For any rational function $\frac{P}{Q}\in\mathbb{F}_{q}(T)$, where $P,Q\in\mathbb{F}_{q}[T]$ and $Q\neq 0$, we set
$\left|\frac{P}{Q}\right|=\begin{cases}q^{\deg(P)-\deg(Q)},&\text{if }P\neq 0,\\ 0,&\text{if }P=0.\end{cases}$
The completion of $\mathbb{F}_{q}(T)$ with respect to this absolute value is the field $\mathbb{F}_{q}((T^{-1}))$, i.e., the field of Laurent series in $T^{-1}$ over the finite field $\mathbb{F}_{q}.$ The absolute value on $\mathbb{F}_{q}((T^{-1}))$, which we again denote by $|\cdot|$, is defined as follows. For $a\in\mathbb{F}_{q}((T^{-1}))$, if $a=0$, then $|a|=0$. Otherwise, $a$ can be expressed uniquely as a Laurent series,
$a=\sum_{k\leq k_{0}}a_{k}T^{k},$
where $k_{0}\in\mathbb{Z}$, each $a_{k}\in\mathbb{F}_{q}$, and the leading coefficient $a_{k_{0}}\neq 0$. We define the degree of $a$ by $\deg(a):=k_{0}$, and set $|a|:=q^{\deg(a)}$. This clearly extends the absolute value $|\cdot|$ on $\mathbb{F}_{q}(T)$ to $\mathbb{F}_{q}((T^{-1}))$, and the extension remains non-archimedean and discrete. Consequently, $\mathbb{F}_{q}((T^{-1}))$ is a complete and separable metric space, which is ultrametric and, hence, totally disconnected. It is worth noting that any local field of positive characteristic is isomorphic to some $\mathbb{F}_{q}((T^{-1}))$.
Throughout the paper, we let $\mathcal{R}$ and $\mathcal{K}$ denote, respectively, the polynomial ring $\mathbb{F}_{q}[T]$ and the field of Laurent series $\mathbb{F}_{q}((T^{-1})).$ We denote by $\mathcal{O}_{\mathcal{K}}$, the ring of integers of $\mathcal{K}$ , defined by $\mathcal{O}_{\mathcal{K}}:=\{x\in\mathcal{K}:|x|\leq 1\}$. Let $\mathcal{L}_{1}$ denote the Haar measure on $\mathcal{K}$, normalized so that $\mathcal{L}_{1}\left(\mathcal{O}_{\mathcal{K}}\right)=1$.
Note that since the root of the equation $x^{\ell}=T$ is not in $\mathcal{K}$ for every $\ell\geq 2$, the degree of the algebraic closure of $\mathcal{K}$ over $\mathcal{K}$ is $\infty$. Nevertheless, one can discuss finite extensions, such as the extension of the function field $\mathcal{K}$ obtained by adjoining the $\ell$-th root of $T$, and we denote this extension by $\mathcal{K}_{\ell}$. Let the polynomial ring in $\mathcal{K}_{\ell}$ be denoted by $\widetilde{\mathcal{R}}$, i.e., $\widetilde{\mathcal{R}}:=\mathbb{F}_{q}[T^{1/\ell}]\subset\mathcal{K}_{\ell}$. So we have the natural inclusion $\mathcal{R}\subset\widetilde{\mathcal{R}}\subset\mathcal{K}_{\ell}.$ It is worth noting that $\widetilde{\mathcal{R}}$ is a lattice in $\mathcal{K}_{\ell}$, while $\mathcal{R}$ is only a discrete subgroup when viewed as a subset of $\mathcal{K}_{\ell}$. Given $n\geq 2$, we equip $\mathcal{K}^{n}$ with the sup norm defined by
$\|\mathbf{x}\|=\max_{1\leq i\leq n}|x_{i}|,\quad\text{for}\,\,\mathbf{x}=(x_{1},\dots,x_{n})\in\mathcal{K}^{n}.$
We also equip $\mathcal{K}_{\ell}^{n}$ with an analogous sup norm. Throughout, $\mathcal{L}_{n}$ denotes the product Haar measure on $\mathcal{K}^{n}$. From now on, we always assume $n:=d+m,$ where $d,m\in\mathbb{N}$.
Given a non-increasing function $\psi:(0,\infty)\to(0,1)$ and a vector $\bm{\Theta}\in\mathcal{K}^{n}$, we define the set of all $(\psi,\bm{\Theta})$-approximable vectors in $\mathcal{K}^{n}$ by
$\mathcal{S}_{n}^{\bm{\Theta}}(\psi):=\left\{\mathbf{x}\in\mathcal{K}^{n}:\left\|\mathbf{x}-\frac{\mathbf{P}+\bm{\Theta}}{Q}\right\|<\frac{\psi(|Q|)}{|Q|}\quad\text{for infinitely many }(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})\right\}.$
In the special case when $\bm{\Theta}=\mathbf{0}$, the set reduces to the set of $\psi$-approximable vectors, which we denote by $\mathcal{S}_{n}({\psi}):=\mathcal{S}_{n}^{\mathbf{0}}(\psi).$
Without loss of generality, we consider the manifold $\mathscr{M}$ defined by the map $\mathbf{f}:U\subset\mathcal{K}^{d}\to\mathcal{K}^{n}$, where
$\mathbf{f}:=(x_{1},\ldots,x_{d},f_{1}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))=(\mathbf{x},\bm{f}(\mathbf{x})),$
with $d=\dim(\mathscr{M})$, $m=\operatorname{\text{codim}}(\mathscr{M})$, and $U\subseteq\mathcal{K}^{d}$ is an open set. We also assume that the defining map $\mathbf{f}$ is analytic (see section˜2 for the definition).
1.2. Main Results
Throughout the paper, we make the following assumptions:
  1. (I)
    The map $\bm{f}=(f_{1},\dots,f_{m}):U\subseteq\mathcal{K}^{d}\rightarrow\mathcal{K}^{m}$ is an analytic map that can be analytically extended to the boundary of $U$.
  2. (II)
    The restrictions of the functions $1,f_{1},\dots,f_{m}$ to any open subset of $U$ are linearly independent over $\mathcal{K}.$
  3. (III)
    We always assume $\psi:(0,\infty)\rightarrow(0,1)$ to be an approximating function, i.e., $\psi$ is non-increasing and $\psi(x)\rightarrow 0$ as $x\rightarrow\infty.$ Furthermore, without loss of generality, we can always assume that the range of $\psi$ is contained in $\{q^{-k}:k\in\mathbb{N}\}.$
Without loss of generality, for the sake of simplification, we further make some boundedness assumptions:
  1. (IV)
    $\|\bm{f}(\mathbf{x})\|\leq 1,$ $\|\nabla\bm{f}(\mathbf{x})\|\leq 1,$ and $|\bar{\Phi}_{\beta}\bm{f}(\mathbf{y_{1}},\mathbf{y_{2}},\mathbf{y_{3}})|\leq M$ for any second-order difference quotient $\Phi_{\beta}$ and $\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},\mathbf{y_{3}}\in U$ (see section˜2 for the definitions).
We now state the main result of this article.
Theorem 1.2.
Let $n\geq 2$, let $\bm{\Theta}\in\mathcal{K}^{n}$, and let $\mathscr{M}\subset\mathcal{K}^{n}$ be an analytic submanifold satisfying (I), (II), and (IV). Suppose that the approximating function $\psi:(0,\infty)\to(0,1)$ satisfies
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}<\infty.$
Then almost every point of $\mathscr{M}$ is not $(\psi,\bm{\Theta})$-approximable, i.e.,
$\mathcal{L}_{d}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0.$
We now recall the definition of Hausdorff $\sigma$-measure. For $\sigma\geq 0$, the Hausdorff $\sigma$-measure of a set $A\subset\mathcal{K}^{n}$ is defined as
$\mathcal{H}^{\sigma}(A):=\lim_{\rho\to 0^{+}}\inf\left\{\sum_{i=1}^{\infty}\operatorname{diam}(A_{i})^{\sigma}:A\subset\bigcup_{i=1}^{\infty}A_{i}\quad\text{ and }\quad\operatorname{diam}(A_{i})<\rho,\,\,\forall\,\,i\right\},$
where $\operatorname{diam}(A)=\sup\{\|\mathbf{u}-\mathbf{v}\|:\mathbf{u},\mathbf{v}\in A\}$ denotes the diameter of a set. Now we state the Hausdorff measure analogue of Theorem 1.2.
Theorem 1.3.
Let $d+m=n\geq 2$, $\bm{\Theta}\in\mathcal{K}^{n}$ and $\psi:(0,\infty)\rightarrow(0,1)$ be an approximating function. Also let $\mathscr{M}\subset\mathcal{K}^{n}$ be an analytic submanifold satisfying (I), (II), and (IV). Suppose that $\sigma>0$ is such that
$\sum_{t\geq 1}\left(\frac{\psi(q^{t})}{q^{t}}\right)^{\sigma+m}q^{(n+1)t}<\infty,$ (1.2)
and
$\sum_{t\geq 1}\left(\frac{\psi(q^{t})}{q^{t}}\right)^{\frac{\sigma-d}{2}}<\infty.$ (1.3)
Then,
$\mathcal{H}^{\sigma}\left(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi))\right)=0.$
Given integers $t,s>0$ and $\Delta\subseteq\mathcal{K}^{n}$, we define the following set, which counts the number of rational points $\frac{{\mathbf{P}}}{Q}$ that are within a distance $q^{-(s+t)}$ of $\mathbf{f}(\Delta\cap U)$,
$\mathcal{R}_{\bm{\Theta}}(\Delta,s,t):=$ $\left\{(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\}):|Q|=q^{t}\ \text{and}\ \inf_{x\in\Delta\cap U}\left\|\mathbf{f}(\mathbf{x})-\frac{{\mathbf{P}}+\bm{\Theta}}{Q}\right\|<q^{-(s+t)}\right\},$
and define $\mathcal{N}_{\bm{\Theta}}(\Delta;s,t)=\#\mathcal{R}_{\bm{\Theta}}(\Delta,s,t)$.
To prove Theorems 1.2 and 1.3, we divide the manifold into two parts: we count the number of rational points near the manifold in the generic part ($U\setminus\mathfrak{M}(s,t)$) and show that the measure of the special part ($\mathfrak{M}(s,t)$) of the manifold is relatively small. For the definition of $\mathfrak{M}(s,t)$, we refer the reader to subsection 4.2. In fact, we prove the following proposition.
Proposition 1.
Let $U\subseteq\mathcal{K}^{d}$ be an open set, and let $\mathbf{f}:U\to\mathcal{K}^{n}$ be a map satisfying (I), (II), and (IV). Then, for any $s,t>0$, there exists a set $\mathfrak{M}(s,t)\subseteq U$ that can be written as the union of disjoint balls in $U$ of radius $q^{-\frac{s+t}{2}}$. For every $\mathbf{x}_{0}\in U$, there exists a ball $B_{0}$ centered at $\mathbf{x}_{0}$ and constants $C,\alpha$ (depending on $B_{0},\mathbf{f}$) such that for every sufficiently large $t>0$, we have
$\mathcal{L}_{d}(\mathfrak{M}(s,t)\cap B_{0})\leq C\max\left\{q^{-\alpha t},q^{\alpha\left(\left(\frac{2n-1}{2(n+1)}\right)s-\frac{3}{2(n+1)}t\right)}\right\}\mathcal{L}_{d}(B_{0}),$
and for every ball $B\subseteq U$ and for every sufficiently large $t$, we have
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t),s,t)\ll_{n,\mathbf{f}}q^{t}\frac{q^{dt}}{q^{ms}}\mathcal{L}_{d}(B),$
where $A\ll_{n,\mathbf{f}}B$ means there exists a constant $c>0$ depending on $n,\mathbf{f}$ such that $A\leq cB.$
1.3. Counting Heuristics
We now motivate the heuristic count leading to Proposition 1. Let $Q\in\mathcal{R}$ be a polynomial with $\text{deg}\ Q=t$. Then, the number of such polynomials $Q$ is
$\{Q\in\mathcal{R}:\deg Q=t\}=(q-1)q^{t-1}\asymp q^{t}.$
For a fixed denominator $Q$ and a fixed $i\in\{1,\dots,n\}$, the number of $P_{i}\in\mathcal{R}$ with $\deg(P_{i})\leq t$ is $q^{t+1}$. Therefore,
$\{{\mathbf{P}}=(P_{1},\dots,P_{n})\in\mathcal{R}^{n}:\deg P_{i}\leq t\}=q^{n(t+1)}.$
Heuristically, this is simplified as follows.
$q^{n(t+1)}=q^{n}\cdot q^{nt}\ll q^{nt},$
since $q^{n}$ is a constant independent of $t$. Hence, the total number of rational points of the form $\frac{{\mathbf{P}}}{Q}\in\mathbb{F}_{q}(T)^{n}$ with $\deg Q=t$, is
$\ll q^{t}\cdot q^{nt}=q^{(n+1)t}.$
Hence, outside a set of small measure, that is the set $\mathfrak{M}(s,t)$ from Proposition 1, the number of rational points $\frac{{\mathbf{P}}}{Q}$ that are within distance $q^{-(s+t)}$ of $\mathbf{f}(\Delta\cap U)$ is
$\ll(q^{-(s+t)})^{m}q^{(n+1)t}=q^{-ms}q^{(d+1)t}.$
1.4. Constraints and Differences in the Positive Characteristic Setting
Although our methodology largely follows the general framework introduced in [BY23], the function field setting poses new difficulties which require a fresh perspective. In the real case, the simultaneous form of Khintchine’s theorem established in [BY23] relies on diagonal flows whose entries are fractional powers of the form $e^{t/\ell}$. The natural analogue in the function field context would involve powers of $T^{\frac{1}{\ell}}$, which do not belong to the base field $\mathcal{K}$, but rather to its extension $\mathcal{K}(T^{\frac{1}{\ell}})$.
To adapt the approach of [BY23], we therefore work over the extended field $\mathcal{K}(T^{\frac{1}{\ell}})$. This, however, introduces additional challenges. In particular, the polynomial ring $\mathcal{R}$ is no longer a lattice in the extension $\mathcal{K}(T^{\frac{1}{\ell}})$, but merely a discrete subgroup. As a result, the classical tools of the geometry of numbers are not directly applicable. To overcome this obstacle, we establish results in the geometry of numbers for specific discrete subgroups, which are not necessarily lattices (see section˜3). Moreover, in the appendix section˜6, we establish a result concerning the first successive minima of such discrete subgroups over $\mathcal{K}(T^{\frac{1}{\ell}})$, which we believe to be of independent interest.
1.5. Structure of the Paper
We begin in section˜2 by recalling the definitions of analytic and $C^{k}$ functions in the ultrametric setting. In section˜3, we establish several auxiliary results in the geometry of numbers over function fields. Proposition 1 is proved in §4 by deriving a measure estimate for the special part of the manifold (see section˜4.3) and a counting estimate for the generic part (see section˜4.2). Finally, Proposition 1 serves as the key ingredient in the proofs of Theorems 1.2 and 1.3, which are presented in section˜5. We conclude the paper with an additional result on the geometry of numbers over function fields, proved in §6.
Acknowledgements
The authors express their sincere gratitude to the University of York for hosting the workshop "Diophantine Approximation and Related Fields," during which a part of this work was done. The authors thank Prof. Anish Ghosh for his continuous encouragement. The fourth-named author would like to thank Prof. Ravi Raghunathan, Prof. Keshav Aggarwal, and Prof. Dipendra Prasad for their interest in this project. The fourth-named author is supported by the Prime Minister’s Research Fellowship (ID: 1302639) during the course of the work.
2. Ultrametric Calculus
In this section, we recall the concept of ultrametric $C^{k}$ functions from [Sch84]. Let $U$ be an open subset of $\mathcal{K}$, and $g:U\subseteq\mathcal{K}\to\mathcal{K}$ be a function. The first-order difference quotient of $g$, denoted by $\Phi^{1}g$, is defined as
$\Phi^{1}g(x,y):=\frac{g(x)-g(y)}{x-y},\quad{for}\,\,(x,y)\in\nabla^{2}U,$
where
$\nabla^{2}U:=\{(x,y)\in U\times U\mid x\neq y\}.$
We say that $g$ is $C^{1}$ at a point $a\in U$ if the limit
$\lim_{(x,y)\to(a,a)}\Phi^{1}g(x,y)$
exists and $g$ is called $C^{1}$ on $U$ if it is $C^{1}$ at every point of $U$. To define a $C^{k}$ function, consider
$\nabla^{k}U:=\{(x_{1},\dots,x_{k})\in U^{k}\mid x_{i}\neq x_{j}\text{ for }i\neq j\},$
and define the $k$-th order difference quotient $\Phi^{k}g:\nabla^{k+1}U\rightarrow\mathcal{K}$ inductively:
$\Phi^{0}g:=g,\quad\Phi^{k}g(x_{1},\dots,x_{k+1}):=\frac{\Phi^{k-1}g(x_{1},x_{3},\dots,x_{k+1})-\Phi^{k-1}g(x_{2},x_{3},\dots,x_{k+1})}{x_{1}-x_{2}}.$
One can easily see that $\Phi^{k}g$ is a symmetric function of its $k+1$ variables. $g$ is said to be $C^{k}$ at a point $a\in U$ if the following limit exists:
$\lim_{(x_{1},\dots,x_{k+1})\to(a,\dots,a)}\Phi^{k}g(x_{1},\dots,x_{k+1})$
and $g$ is said to be $C^{k}$ on $U$, i.e., $g\in C^{k}(U)$ if $g$ is $C^{k}$ at every point of $U$. This is equivalent to $\Phi^{k}g$ admitting an extension $\bar{\Phi}^{k}g:U^{k+1}\rightarrow{\mathcal{K}}$, which, if it exists, is unique. If $g\in C^{k}(U)$, all derivatives up to order $k$ exist and the $k$-th derivative of $g$ is given by:
$g^{(k)}(x)=k!\,\bar{\Phi}^{k}g(x,\dots,x).$
Now let $g:U_{1}\times\dots\times U_{d}\subseteq\mathcal{K}^{d}\rightarrow\mathcal{K}$ be a $\mathcal{K}$-valued function of several variables, where $U_{i}\subseteq\mathcal{K}$ are open subsets for $i=1,\dots,d$. Let $\Phi_{i}^{k}g$ denote the $k$-th order difference quotient with respect to the $i$-th variable. For a multi-index $\beta=(i_{1},\dots,i_{d})$, we define:
$\Phi_{\beta}g:=\Phi_{1}^{i_{1}}\circ\cdots\circ\Phi_{d}^{i_{d}}g.$
The domain of $\Phi_{\beta}g$ is $\nabla^{i_{1}+1}U_{1}\times\cdots\times\nabla^{i_{d}+1}U_{d}.$ We say $g\in C^{k}(U_{1}\times\cdots\times U_{d})$ if for all multiindex $\beta=(i_{1},\dots,i_{d})$ with $|\beta|:=\sum_{j=1}^{d}i_{j}\leq k$, the difference quotient $\Phi_{\beta}g$ extends continuously to
$\bar{\Phi}_{\beta}g:U_{1}^{i_{1}+1}\times\cdots\times U_{d}^{i_{d}+1}\to\mathcal{K}.$
As in the one-variable case, for $g\in C^{k}(U_{1}\times\dots\times U_{d})$, the mixed partial derivatives
$\partial_{\beta}g:=\partial_{1}^{i_{1}}\circ\cdots\circ\partial_{d}^{i_{d}}g$
exist and are continuous for all $|\beta|\leq k$. They also satisfy:
$\partial_{\beta}g(x_{1},\dots,x_{d})=\beta!\,\bar{\Phi}_{\beta}g(x_{1},\dots,x_{1},\dots,x_{d},\dots,x_{d}),$ (2.1)
where each variable $x_{j}$ appears $i_{j}+1$ times, and $\beta!:=\prod_{j=1}^{d}i_{j}!$. Now we define an analytic function in $\mathcal{K}.$
Definition 2.1.
Let $U$ be an open subset of $\mathcal{K}$. A function $g:U\subseteq\mathcal{K}\to\mathcal{K}$ is called analytic on $U$ if for every point $x_{0}\in U$, there exists an open ball
$B(x_{0},r)=\{x\in\mathcal{K}:|x-x_{0}|<r\}\subseteq U$
and a power series
$g(x)=\sum_{n=0}^{\infty}a_{n}(x-x_{0})^{n}$
with coefficients $a_{n}\in\mathcal{K}$, such that this series converges for all $x\in B(x_{0},r)$, and equals $g(x)$ on that ball. In other words, $g$ is analytic on $U$ if it is locally given by a convergent power series at every point in $U$.
Given an analytic map $g=(g_{1},\dots,g_{n}):U\subseteq\mathcal{K}^{d}\to\mathcal{K}^{n}$, by $\nabla g(\mathbf{x})$ we denote the $d\times n$ matrix whose $(i,j)$-th entry is $\partial_{j}g_{i}(x)$. We also recall the following second-order Taylor expansion formula for analytic functions (see [BDG24, §11-Appendix] for details).
Lemma 2.2.
Let $U$ be an open subset of $\mathcal{K}^{d}$ and $\mathbf{x},\mathbf{x}^{\prime}\in U$. Also, let $g:U\to\mathcal{K}^{n}$ be an analytic function. Then, we have
$g(\mathbf{x}+\mathbf{x}^{\prime})=g(\mathbf{x})+\sum_{i=1}^{d}x_{i}^{\prime}\bar{\Phi}_{e_{i}}g(\cdot)+\sum_{\beta=(i_{1},\ldots,i_{d}),|\beta|=2}\bar{\Phi}_{\beta}g(\cdot)\prod_{i=1}^{d}(x_{i}^{\prime})^{i_{j}},$
where $e_{i}$ is the multiindex whose $i$-th coordinate is $1$ and all other coordinates are zero, and the arguments of $\bar{\Phi}_{e_{i}}g(\cdot)$ and $\bar{\Phi}_{\beta}g(\cdot)$ are some of the components of $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
In particular, in our context, applying the above formula for $f_{i}$’s and using the bounds of the first and second difference quotients of $f_{i}$ (see (IV)), we have
$\|f_{i}(\mathbf{x}+\mathbf{x}^{\prime})-f_{i}(\mathbf{x})\|\leq\max\{M\|\mathbf{x}^{\prime}\|,M\|\mathbf{x}^{\prime}\|^{2}\}\,\,\,\,\text{for}\,\,i=1,\dots,m.$ (2.2)
We conclude this section with the following version of Besicovitch’s covering theorem, which will be used in section˜4.
Theorem 2.3.
[KT07, Example 2.1] Let $A$ be a bounded subset of $\mathcal{K}^{d}$. Then any covering by balls of $A$ has a countable subcover consisting of mutually disjoint balls.
3. Preliminaries in Geometry of Numbers
Recall that for $\ell\in\mathbb{N}$, $\mathcal{K}_{\ell}=\mathcal{K}(T^{1/\ell})$ and $\mathcal{R}=\mathbb{F}_{q}[T]$. Note that $\mathcal{K}_{\ell}=\mathcal{K}$ when $\ell=1$. For $i=1,\dots,n+1$ and $g\in\operatorname{SL}(n+1,\mathcal{K}_{\ell})$, we define the $i$-th successive minimum of the discrete subgroup $g\mathcal{R}^{n+1}\subseteq\mathcal{K}_{\ell}^{n+1}$ as
$\lambda_{i}(g\mathcal{R}^{n+1}):=\inf\left\{\lambda>0:B(0,\lambda)\cap g\mathcal{R}^{n+1}\text{ contains }i\text{ linearly independent vectors over }\mathcal{K}_{\ell}\right\},$
where $B(0,\lambda)$ denotes the ball of radius $\lambda$ centered at the origin in $\mathcal{K}_{\ell}^{n+1}$.
We define $G:=\operatorname{SL}(n+1,\mathcal{K})$ and $\Gamma:=\operatorname{SL}(n+1,\mathcal{R})$. Then the associated homogeneous space $\mathcal{X}_{n+1}:=G/\Gamma$ can be identified with the space of all unimodular lattices in $\mathcal{K}^{n+1}$. Now we recall the function field analogue of Minkowski’s second theorem from [Mah41, Equations (24) and (25)].
Theorem 3.1.
Let $g\in\operatorname{GL}(n+1,\mathcal{K})$. Then,
$\prod_{i=1}^{n+1}\lambda_{i}(g\mathcal{R}^{n+1})=|\det(g)|.$ (3.1)
Given a lattice $\Lambda\in\mathcal{X}_{n+1}$, we now define the dual lattice of $\Lambda$ as follows
$\Lambda^{*}:=\left\{\mathbf{y}\in\mathcal{K}^{n+1}:\forall\mathbf{x}\in\Lambda,\langle\mathbf{x},\mathbf{y}\rangle\in\mathcal{R}\right\},$
where $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We first state the duality theorem on $\mathcal{K}^{n+1}$, originally due to Mahler[Mah41, Equation (28)] (see also [BK25, Lemma 5.4]).
Theorem 3.2.
Let $\Lambda$ be a lattice in $\mathcal{K}^{n+1}$, and $\Lambda^{*}$ be its dual lattice. Then we have
$\lambda_{i}(\Lambda^{*})\lambda_{n+2-i}(\Lambda)=1,\ \text{for}\ 1\leq i\leq n+1.$ (3.2)
For integers $a_{1},\dots,a_{n+1}$, consider the matrix
$g=\operatorname{diag}\left(T^{\frac{a_{1}}{\ell}},\dots,T^{\frac{a_{n+1}}{\ell}}\right),$ (3.3)
and $g^{*}$ is defined as $g^{*}:=(g^{t})^{-1}$. We prove the following generalization of Theorem 3.2 for some specific discrete subgroups in $\mathcal{K}_{\ell}^{n+1}$.
Theorem 3.3 (Duality).
Let $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$ be of the form (3.3) and $\Lambda$ be a lattice in $\mathcal{K}^{n+1}$. Then
$q^{-\frac{\ell-1}{\ell}}\leq\lambda_{i}(g^{*}\Lambda^{*})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}},~\forall i=1,\dots,n+1.$ (3.4)
Proof.
Without any loss of generality, we may assume that $a_{i}\in\left\{0,1,\dots,\ell-1\right\}$; for, if necessary, we may replace $\Lambda$ by the lattice $\operatorname{diag}\left(T^{\left\lfloor\frac{a_{1}}{\ell}\right\rfloor},\dots,T^{\left\lfloor{\frac{a_{n+1}}{\ell}}\right\rfloor}\right)\Lambda$, where $\lfloor c\rfloor:=\max\{k\in\mathbb{Z}:k\leq c\}$ for $c\in\mathbb{R}.$ Note that,
$\|\mathbf{v}\|=\|g^{-1}g\mathbf{v}\|\leq\|g^{-1}\|\|g\mathbf{v}\|\leq\|g\mathbf{v}\|\leq\|g\|\|\mathbf{v}\|\leq q^{\frac{\ell-1}{\ell}}\|\mathbf{v}\|;\text{ and }$
$q^{-\frac{\ell-1}{\ell}}\|\mathbf{v}\|\leq\|g\|^{-1}\|\mathbf{v}\|=\|g\|^{-1}\|gg^{-1}\mathbf{v}\|\leq\|g^{-1}\mathbf{v}\|\leq\|\mathbf{v}\|,\forall\mathbf{v}\in\mathcal{K}^{n+1}.$ (3.5)
Since $\mathcal{K}_{\ell}$ is a finite extension over $\mathcal{K}$, then $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}\in\Lambda$ are linearly independent over $\mathcal{K}$ if and only if $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}$ are linearly independent over $\mathcal{K}_{\ell}$. Hence,
[…]
Consider linearly independent vectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}$ in $\Lambda$ such that $\|\mathbf{v}_{i}\|=\lambda_{i}(\Lambda)=\lambda_{n+2-i}^{-1}\left(\Lambda^{\ast}\right)$, for all $i=1,\dots,n+1$. Then from (3.6) and (3.5), it follows that
$\lambda_{i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda),~\forall i=1,\dots,n+1.$ (3.7)
Similarly, one obtains that
$\lambda_{i}(\Lambda)\leq\lambda_{i}(g\Lambda),$ (3.8)
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda)\leq\lambda_{i}(g^{-1}\Lambda),\text{ and }$ (3.9)
$\lambda_{i}(g^{-1}\Lambda)\leq\lambda_{i}(\Lambda),~\forall i=1,\dots,n+1.$ (3.10)
Since $\Lambda$ is arbitrary, by replacing $\Lambda$ with $\Lambda^{*}$, from (3.9) and (3.10), we also have
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\leq\lambda_{i}(\Lambda^{\ast}),~\forall i=1,\dots,n+1,$ (3.11)
since $g$ is diagonal, so that $g^{-1}$ is the same as its transpose. Multiplying inequalities in (3.7) with j=n+2ij=n+2-i and (3.11) now yields that
$\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda),~\forall i=1,\dots,n+1.$
On the other hand, using (3.8) and (3.11), we obtain,
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda)\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda).$
Combining these, we have,
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda)\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda),~\forall i=1,\dots,n+1.$
Equation (3.4) is immediate from this, due to (3.2). ∎
Remark 3.4.
When $\ell=1$, (3.4) reduces to (3.2).
Recall from [Mah41] that convex bodies in $\mathcal{K}^{n+1}$ are precisely the sets of the form $g\mathcal{O}_{\mathcal{K}}^{n+1}$, where $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$. Then, clearly $\operatorname{Vol}(g\mathcal{O}_{\mathcal{K}}^{n+1})=|\det(g)|$. We now state the following lemma, due to Bagshaw and Kerr [BK25], that counts the number of lattice points in a convex body:
Lemma 3.5.
[BK25, Lemma 6.2] Let $u\in\operatorname{GL}(n+1,\mathcal{K})$, let $\Lambda=u{\mathcal{R}}^{n+1}$, and let $\mathcal{C}=h\mathcal{O}_{\mathcal{K}}^{n+1}$ be a convex body. Then,
$(\Lambda\cap\mathcal{C})=\prod_{i=1}^{n+1}\left\lceil\frac{q}{\lambda_{i}(h^{-1}\Lambda)}\right\rceil,$
where $\lceil\cdot\rceil:\mathbb{R}\rightarrow\mathbb{Z}$ is defined by $\lceil\alpha\rceil=\min\{k\in\mathbb{Z}:k\geq\alpha\}$. In particular, if $\mathcal{C}$ contains a fundamental domain for $\Lambda$, then,
$(\Lambda\cap\mathcal{C})=q^{n+1}\frac{\operatorname{Vol}(\mathcal{C})}{|\det(u)|}.$
We generalize the above result for specific discrete subgroups of the form $g\Lambda\subseteq\mathcal{K}_{\ell}^{n+1}$ with $g$ being diagonal over $\mathcal{K}_{\ell}$ and $\Lambda$ being a lattice in $\mathcal{K}$ under the assumption $\lambda_{n+1}(g\Lambda)\leq q^{c}$ for some $c>0$.
Lemma 3.6.
Let $g=\operatorname{diag}\left(T^{a_{1}+\frac{\alpha_{1}}{\ell}},\dots,T^{a_{n+1}+\frac{\alpha_{n+1}}{\ell}}\right)\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$, where $a_{i}\in\mathbb{Z}$ and $\alpha_{i}\in\{0,1,\dots,(\ell-1)\}$. Consider any $c\geq 0$ such that $\lambda_{n+1}(g\Lambda)\leq q^{c}$, where $\Lambda=u\mathcal{R}^{n+1}$ and $u\in\operatorname{GL}(n+1,\mathcal{K})$. Then,
$\left\{\mathbf{v}\in g\Lambda:\|\mathbf{v}\|\leq q^{c}\right\}\leq\frac{q^{(n+1)(c+1)}}{|\det(g)|\cdot|\det(u)|}.$
Proof.
Denote $g]=\operatorname{diag}(T^{a_{1}},\dots,T^{a_{n+1}})$ and $\{g\}=\operatorname{diag}\left(T^{\frac{\alpha_{1}}{\ell}},\dots,T^{\frac{\alpha_{n+1}}{\ell}}\right)$. Since $g=\{g\}[g]$ and $|\det(\{g\})|\geq 1$, clearly $|\det(g)|\geq|\det([g])|$.
$\|[g]\mathbf{v}\|=\|\{g\}^{-1}{g}\mathbf{v}\|\leq\|{g}\mathbf{v}\|,\text{ for all }\mathbf{v}\in\Lambda.$ (3.12)
Let $g\mathbf{v}_{1},\dots,g\mathbf{v}_{n+1}$ be linearly independent vectors in $g\Lambda$ with $\|g\mathbf{v}_{i}\|=\lambda_{i}(g\Lambda)$. Clearly, $g]\mathbf{v}_{1},\dots,[g]\mathbf{v}_{n+1}$ are linearly independent. Then it follows from (3.12) and the hypothesis $\lambda_{n+1}(g\Lambda)\leq q^{c}$ that $\lambda_{n+1}([g]\Lambda)\leq q^{c}$. Furthermore, (3.12) yields that the parallelepiped spanned by the vectors $g]\mathbf{v}_{1},\dots,[g]\mathbf{v}_{n+1}$ inside $\mathcal{K}^{n+1}$ is actually contained in $T^{\lfloor{c}\rfloor}\mathcal{O}_{\mathcal{K}}^{n+1}$, where $\lfloor c\rfloor=\max\{n\in\mathbb{Z}:n\leq c\}$. Now, in view of (3.12), applying Lemma 3.5, we arrive at the following:
$\begin{split}\#\left\{\mathbf{v}\in\{g\}[g]\Lambda:\|\mathbf{v}\|\leq q^{c}\right\}\leq\#\{\mathbf{v}\in[g]\Lambda:\|\mathbf{v}\|\leq q^{\lfloor{c}\rfloor}\}\leq\frac{q^{(n+1)(c+1)}}{|\det([g])|\cdot|\det(u)|}\leq\frac{q^{(n+1)(c+1)}}{|\det(g)|\cdot|\det(u)|}.\end{split}$
4. Analysis on generic and special part
4.1. Some Preliminary Estimates
In this subsection, we provide several estimates that will be used to prove Theorems 1.2 and 1.3. Recall that the manifold $\mathscr{M}$ is defined by the map $\mathbf{f}:U\subseteq\mathcal{K}^{d}\to\mathcal{K}^{n}$, where
$\mathbf{f}:=(x_{1},\ldots,x_{d},f_{1}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))=(\mathbf{x},\bm{f}(\mathbf{x})),\,\,\text{for}\,\,\mathbf{x}=(x_{1},\dots,x_{d})\in U$
with $d=\dim(\mathscr{M})$, $m=\operatorname{\text{codim}}(\mathscr{M})$, and $U\subseteq\mathcal{K}^{d}$ being an open set. We also assume that the defining map $\mathbf{f}$ is analytic.
Lemma 4.1.
Let $s,t\in\mathbb{N}$ and suppose that for some $\mathbf{x}\in U$, we have
$\left\lVert\mathbf{f}(\mathbf{x})-\frac{\mathbf{P}+\bm{\Theta}}{Q}\right\rVert<q^{-(s+t)},$
where $\mathbf{P}=(P_{1},\cdots,P_{n})\in\mathcal{R}^{n},$ $Q\in\mathcal{R}\setminus\{0\}$ with $|Q|=q^{t}$ and $\bm{\Theta}=(\Theta_{1},\cdots,\Theta_{n})$. Then,
  1. (1)
    $|Qx_{i}-P_{i}-\Theta_{i}|<q^{-s}$ for every $i=1,\dots,d$.
  2. (2)
    $|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(Qx_{i}-P_{i}+\Theta_{i})-P_{d+j}+\Theta_{d+j}|<\max\{1,M\}q^{-s}$ for every $j=1,\dots,m$.
Proof.
(1) is easily seen by considering the $i$-th coordinate of $\mathbf{f}(\mathbf{x})-\mathbf{P}-\bm{\Theta}$ for $i=1,\dots,d$, and using $|Q|=q^{t}$. To prove (2), using the ultrametric inequality, we obtain
$\bigg|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})\big(Qx_{i}-P_{i}-\Theta_{i}\big)-P_{d+j}-\Theta_{d+j}\bigg|$ $\leq\max\Big\{|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|,$
$\max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Big\}.$
Now, from the conditions $\big\|\mathbf{f}(\mathbf{x})-\tfrac{\mathbf{P}+\bm{\Theta}}{Q}\big\|<q^{-(s+t)}\,\,\text{and}\,\,|Q|=q^{t}$, it follows that
$|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|<q^{-s}.$
Combining this and (1), we conclude
$\max\Big\{|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|,\max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Big\}<\max\{1,M\}q^{-s}.$
Lemma 4.2.
Suppose that $\mathbf{x}\in U$ and $(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})$ satisfy the hypothesis of Lemma 4.1. Let $\mathbf{x}\in B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})$ for some $\mathbf{x}_{0}=(x_{1,0},\cdots,x_{d,0})\in U$, then,
  1. (1)
    $|Qx_{i,0}-P_{i}-\Theta_{i}|<\max\{q^{-s},q^{t-\frac{s+t}{2}}\}$ for $i=1,\dots,d$.
  2. (2)
    For every $j=1,\dots,m$, we have
    $\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{0,i}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|\leq\max\{1,M\}q^{-s}.$
Proof.
Let $\mathbf{x}_{0}=(x_{1,0},\cdots,x_{d,0})\in U$. By Lemma 4.1, we have the vector $\mathbf{x}$ satisfies (1) and (2). Hence,
$|Qx_{i,0}-P_{i}-\Theta_{i}|$ $\leq$ $\max\left\{|Qx_{i}-P_{i}-\Theta_{i}|,|Q|\cdot|x_{i}-x_{i,0}|\right\}$
$\leq$ $\max\{q^{-s},q^{t-\frac{s+t}{2}}\}.$
Note that by Lemma 2.2, we have
$\begin{split}\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{0,i}-P_{i}+\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|\\ \leq\max\Bigg\{\left|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(Qx_{i}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|,\\ |Q|\cdot\left|f_{j}(\mathbf{x}_{0})-f_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(x_{0,i}-x_{i})\right|,\\ \max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x}_{0})-\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Bigg\}\\ \leq\max\{1,M\}q^{-s}.\end{split}$
4.2. Analysis on Generic Part
We first define the generic part. To do so, we first define the diagonal flow $g_{s,t}$ as
$g_{s,t}:=\operatorname{diag}\left\{\underbrace{T^{\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{\frac{(d+2)t+2s}{2(n+1)}}}_{m},\overbrace{T^{\frac{(d+2)t+2s}{2(n+1)}-\frac{s+t}{2}},\dots,T^{\frac{(d+2)t+2s}{2(n+1)}-\frac{s+t}{2}}}^{d},T^{\frac{(d+2)t+2s}{2(n+1)}-s-t}\right\}.$
Note that the entries of the diagonal flow $g_{s,t}$ come from the extended field $\mathcal{K}_{2(n+1)}$. For $\mathbf{x}\in U$, denote by $J(\mathbf{x})$ the Jacobian matrix of $\mathbf{f}$ at $\mathbf{x}$, i.e.,
$J(\mathbf{x}):=\left(\frac{\partial f_{i}}{\partial x_{j}}(\mathbf{x})\right)_{1\leq i\leq d,1\leq j\leq m}.$
Given $k\in\mathbb{N}$, we let $\sigma_{k}$ stand for the following $k\times k$ matrix:
$\sigma_{k}:=\begin{pmatrix}0&0&\ldots&0&1\\ 0&0&\ldots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\ldots&0&0\\ 1&0&\ldots&0&0\end{pmatrix}$.
Observe that $\sigma_{k}$ acts on row vectors from the right and on column vectors from the left by reversing the order of their coordinates. $\sigma_{k}$ is an involution. Clearly $\sigma_{k}^{-1}=\sigma_{k}$. We now also consider the following matrices in $\operatorname{SL}(n+1,\mathcal{K})$:
[…]
It can be seen that $u_{1}(\mathbf{x})$ has the following form:
$u_{1}(\mathbf{x})=\begin{pmatrix}1&0&\dots&0&-\partial_{d}f_{m}(\mathbf{x})&\dots&-\partial_{1}f_{m}(\mathbf{x})&f_{m}(\mathbf{x})-\sum_{i=1}^{d}x_{i}\partial_{i}f_{m}(\mathbf{x})\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\vdots&1&-\partial_{d}f_{1}(\mathbf{x})&\dots&-\partial_{1}f_{1}(\mathbf{x})&f_{1}(\mathbf{x})-\sum_{i=1}^{d}x_{i}\partial_{i}f_{1}(\mathbf{x})\\ 0&0&\dots&0&1&0&\dots&x_{d}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\dots&0&0&\dots&1&x_{1}\\ 0&0&\dots&0&0&\dots&0&1\end{pmatrix}$
We define a new operator on $\operatorname{GL}(n+1,\mathcal{K}_{\ell})$ by
$g^{\star}:=\sigma_{n+1}^{-1}\left(g^{t}\right)^{-1}\sigma_{n+1}\quad\text{for }g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell}).$
Note that given any $g_{1},g_{2}\in\operatorname{GL}(n+1,\mathcal{K}_{\ell}),$ one has $(g_{1}g_{2})^{\star}=g_{1}^{\star}g_{2}^{\star}$. Since $\sigma_{n+1}$ acts by a mere permutation of the coordinates, for any $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$, we have $\lambda_{i}\left(\left(\sigma_{n+1}^{-1}g\sigma_{n+1}\right)\mathcal{R}^{n+1}\right)=\lambda_{i}(g\mathcal{R}^{n+1})$ for every $i=1,\dots,n+1$. By replacing $g$ with $\left(g^{t}\right)^{-1}$, for every $i=1,\dots,n+1$, we have $\lambda_{i}\left(g^{\star}\mathcal{R}^{n+1}\right)=\lambda_{i}\left(g^{*}\mathcal{R}^{n+1}\right)$, where $g^{*}=(g^{t})^{-1}$ is the dual of $g$ as defined in section˜3.
Define
$\begin{split}\mathfrak{M}_{0}(s,t):=\left\{\mathbf{x}\in U:\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\}\\ =\left\{\mathbf{x}\in U:\lambda_{1}(g_{s,t}^{*}u_{1}^{*}(\mathbf{x})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\},\end{split}$ (4.1)
and
$\mathfrak{M}(s,t)=\bigcup_{\mathbf{x}\in\mathfrak{M}_{0}(s,t)}B(\mathbf{x},q^{-\frac{s+t}{2}}).$
Henceforth, we define the generic part of the manifold as $U\setminus\mathfrak{M}(s,t)$ and the special part as $\mathfrak{M}(s,t)$.
We first estimate the number of points in the generic part of the manifold. Towards this end, we prove the following proposition.
Proposition 2.
Let $\mathbf{f}$ and $U$ be as in Proposition 1. Then, for any $s\in\mathbb{N}$, for any ball $B\subset U$ , and all sufficiently large $t$, we have
$\mathcal{N}_{\bm{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\ll q^{(d+1)t}q^{-ms}\mathcal{L}_{d}(B).$ (4.2)
The proof of Proposition 2 relies on the following counting estimate:
Lemma 4.3.
Let $\mathbf{f}$ and $U$ be as in Proposition 1. Also let $B$ be an open ball inside $U\subset\mathcal{K}^{d}$. Then, for all $\mathbf{x}_{0}\in(U\setminus\mathfrak{M}(s,t))\cap B$ and $t>0$ sufficiently large,
$\mathcal{N}_{\bm{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t)\ll q^{\frac{d}{2}(t+s)+(t-ns)}.$
Proof.
Without loss of generality, assume that $\mathcal{N}_{\bm{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t)>1$. Then for $i=1,2$ there exist $(\mathbf{P}_{i},Q_{i})\in\mathcal{R}^{n+1}$ and $\mathbf{x}_{i}\in B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})$, such that
$\left\|\mathbf{f}(\mathbf{x}_{i})-\frac{\mathbf{P}_{i}+\bm{\Theta}}{Q_{i}}\right\|<q^{-(s+t)}.$ (4.3)
Hence, by Lemma 4.2, we have for $h=1,2$, and for every $i=1,\dots,d,$, we have
$|Q_{h}x_{0,i}-P_{h,i}-\Theta_{i}|<\max\{q^{-s},q^{t-\frac{s+t}{2}}\}.$ (4.4)
Now, for every $j=1,\dots,m$, we have
$\begin{split}\left|Q_{h}f_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Q_{h}x_{0,i}-P_{h,i}-\Theta_{i})-P_{h,d+j}-\Theta_{d+j}\right|\leq\max\{1,M\}q^{-s}.\end{split}$ (4.5)
Let $Q=Q_{2}-Q_{1}$ and let $\mathbf{P}=\mathbf{P}_{2}-\mathbf{P}_{1}$, so that $|Q|\leq q^{t}$. Hence, by subtracting (4.4) from each other with $h=1,2$, we have
$|Qx_{0,i}-P_{i}|<\max\left\{q^{-s},q^{t-\frac{s+t}{2}}\right\}=\max\left\{q^{-s},q^{\frac{t-s}{2}}\right\}=q^{\frac{t-s}{2}},$ (4.6)
whenever $t$ is large enough. Similarly, by subtracting (4.5) from each other with $h=1,2$, we have
$\begin{split}\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{0,i}-P_{i})-P_{d+j}\right|\leq\max\{1,M\}q^{-s}.\end{split}$ (4.7)
Given $r>0$ and $k\in\mathbb{N}$, define $r]^{k}=\{\mathbf{v}\in\mathcal{K}_{2(n+1)}^{k}:\|\mathbf{v}\|\leq r\}$. Thus, we have
$u_{1}(\mathbf{x}_{0})\begin{pmatrix}-\mathbf{P}\sigma_{n}\\ Q\end{pmatrix}\in\left(\left[\max\{1,M\}q^{-s}\right]^{m}\times\left[q^{\frac{t-s}{2}}\right]^{d}\times[q^{t}]\right)\cap u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}.$
Thus,
$g_{s,t}u_{1}(\mathbf{x}_{0})\begin{pmatrix}-\mathbf{P}\sigma_{n}\\ Q\end{pmatrix}\in\overbrace{\left(\max\{1,M\}\left(\left[q^{\frac{(d+2)t-2ns}{2(n+1)}}\right]^{n+1}\right)\right)}^{\Omega}\cap g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}.$
Hence, one has $\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x}_{0})\mathcal{R}^{n+1})=\lambda_{1}(g_{s,t}^{*}u_{1}^{*}(\mathbf{x}_{0})\mathcal{R}^{n+1})\geq q^{-\frac{(d+2)t-2ns}{2(n+1)}}.$ Now, it follows from duality (Theorem 3.3) that $\lambda_{n+1}(g_{s,t}u_{1}(\mathbf{x_{0}}))\mathcal{R}^{n+1})\ll_{n}q^{\frac{(d+2)t-2ns}{2(n+1)}}$. Therefore, by applying Lemma 3.6 with $c=\frac{(d+2)t-2ns}{2(n+1)}$, for every sufficiently large $t>0$, we have
$\#g_{s,t}u_{1}(\mathbf{x}_{0})\mathcal{R}^{n+1}\cap\max\{1,M\}\left[q^{\frac{(d+2)t-2ns}{2(n+1)}}\right]^{n+1}$ $\ll\frac{q^{\frac{(d+2)t-2ns}{2}}}{q^{\frac{(d+2)t+2s}{2}-\frac{d+2}{2}(s+t)}}$
$=q^{\left(\frac{d}{2}+1\right)t-s\left(n-\frac{d}{2}\right)}$
$=q^{\frac{d}{2}(t+s)+(t-ns)}.$
The following lemma, coupled with Lemma 4.3, now yields Proposition 2:
Lemma 4.4.
For all sufficiently large $t>0$, we have
$\mathcal{N}_{\mathbf{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\leq q^{\frac{d(t+s)}{2}}\mathcal{L}_{d}(B)\max_{\mathbf{x}_{0}\in B\setminus\mathfrak{M}(s,t)}\mathcal{N}_{\mathbf{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t).$
Proof.
The proof of this Lemma is similar to that of [BY23, Lemma 5.4]. Firstly, one can cover $B$ with at most $q^{\frac{d}{2}(s+t)}\mathcal{L}_{d}(B)$ balls of radius $q^{-\frac{s+t}{2}}$. If one of these balls intersects $\mathfrak{M}(s,t)$, then there exists $\mathbf{x}_{0}\in\mathfrak{M}(s,t)$ such that the ball coincides with $B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})$. Therefore,
$\mathcal{N}_{\mathbf{\Theta}}(B\setminus\mathfrak{M}(s,t);s,t)\leq q^{\frac{d}{2}(t+s)}\mathcal{L}_{d}(B)\max_{\mathbf{x}_{0}\in B\setminus\mathfrak{M}(s,t)}\mathcal{N}_{\mathbf{\Theta}}(B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})\cap B;s,t).$
4.3. Analysis on Special Part
To deal with the special part of the manifold, for $r,s^{\prime},t_{1},\dots,t_{n}\in\mathbb{Z}$ and for any ball $B\subset\mathcal{K}^{d}$, define
$\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})=\left\{\mathbf{x}\in B:\exists\,\,\mathbf{P}=(P_{1},\dots,P_{n})\in\mathcal{R}^{n},Q_{0}\in\mathcal{R}\left|\begin{array}[]{lcl}|\mathbf{f}(\mathbf{x})\cdot\mathbf{P}+Q_{0}|<q^{-r}\\[8.61108pt] \|\nabla(\mathbf{f}(\mathbf{x})\cdot\mathbf{P})\|<q^{s^{\prime}}\\[8.61108pt] |P_{i}|<q^{t_{i}},\,\,\,i=1,\dots,n\end{array}\right.\right\}$.
We use the following theorem by Das and Ganguly [DG22], which bounds the measure of the set $\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})$.
Theorem 4.5 ([DG22, Theorem 6.1]).
Suppose $U$ is an open subset of $\mathcal{K}^{d}$ and $\mathbf{f}:U\rightarrow\mathcal{K}^{n}$ satisfies (II), (III) and (IV). Then for any $\mathbf{x}_{0}\in U,$ one can find a neighbourhood $V\subseteq U$ of $\mathbf{x}_{0}$ and $\alpha>0$ with the following property: for any ball $B\subseteq V,$ there exists $E>0$ such that for any choice of $r,s^{\prime},t_{1},\dots,t_{n}\in\mathbb{Z}$ with $r\geq 0,$ $t_{1},\dots,t_{n}\geq 1,$ and $s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}<0$ one has
$\mathcal{L}_{d}\left(\mathfrak{G}_{\mathbf{f}}(r,s^{\prime},t_{1},\dots,t_{n})\right)\leq E\gamma^{\alpha}\mathcal{L}_{d}(B),$ (4.8)
where $\gamma:=\max\left(q^{-r},q^{\frac{s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}}{n+1}}\right)$.
We now prove that the special part of the manifold $\mathfrak{M}(s,t)$ is contained inside the set $\mathfrak{G}_{f}(\widetilde{C}t,\widetilde{C}\frac{s-t}{2},\widetilde{C}s,\dots,\widetilde{C}s)$ for sufficiently large $t$ and some constant $\widetilde{C}$.
Lemma 4.6.
Let $\mathbf{f}$ and $U$ be as in Proposition 1. For $t$ large enough, there exists a constant $\widetilde{C}\in\mathbb{N}$ (depending on $M,d$ and $n$) such that we have
$\begin{split}\mathfrak{M}(s,t)\subseteq\mathfrak{G}_{\mathbf{f}}\left(\widetilde{C}t,\widetilde{C}\frac{s-t}{2},\widetilde{C}s,\dots,\widetilde{C}s\right).\end{split}$
Proof.
Let $\mathbf{x}_{0}\in\mathfrak{M}_{0}(s,t)$. Then,
$\lambda_{1}(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x}_{0})\mathcal{R}^{n+1})<q^{-\frac{(d+2)t-2ns}{2(n+1)}}.$ (4.9)
Note that
$g_{s,t}^{\star}=\operatorname{diag}\left(T^{s+t-\frac{(d+2)t+2s}{2(n+1)}},T^{\frac{s+t}{2}-\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{\frac{s+t}{2}-\frac{(d+2)t+2s}{2(n+1)}},T^{-\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{-\frac{(d+2)t+2s}{2(n+1)}}\right)$
and $u_{1}^{\star}(\mathbf{x}_{0})=\begin{pmatrix}1&-\mathbf{x}_{0}&-f(\mathbf{x}_{0})\\ &\mathrm{I}_{d}&J(\mathbf{x}_{0})\\ &&\mathrm{I}_{m}\end{pmatrix}$. Then, by (4.9), there exists some $(Q_{0},\mathbf{P}):=(Q_{0},\mathbf{P}^{(d)},\mathbf{P}^{(m)})=(Q_{0},P_{1}^{(d)},\dots,P_{d}^{(d)},P_{1}^{(m)},\dots,P_{m}^{(m)})\in\mathcal{R}^{n+1}\setminus\{0\}$, such that
$\begin{array}[]{lcl}|Q_{0}+\mathbf{P}\cdot\mathbf{f}(\mathbf{x}_{0})|<q^{-t},\\[8.61108pt] \left|{P}^{(d)}_{i}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|<q^{\frac{s-t}{2}},\,\text{ for }i=1,\dots,d\\[8.61108pt] |{P}^{(m)}_{j}|<q^{s},\,\text{ for }j=1,\dots,m.\end{array}$ (4.10)
By the second and third inequalities in (4.10), for every $i=1,\dots,d$,
$\begin{split}\left|{P}^{(d)}_{i}\right|&\leq\max\left\{\left|{P}^{(d)}_{i}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|,\max_{j=1,\dots,m}\left|\partial_{i}f_{j}(\mathbf{x}_{0}){P}^{(m)}_{j}\right|\right\}\\ &\leq\max\{q^{-t},Mq^{s}\}\\ &\leq\max\{1,M\}q^{s}.\end{split}$ (4.11)
If $\mathbf{x}\in\mathfrak{M}(s,t)$, then there exists $\mathbf{x}_{0}\in\mathfrak{M}_{0}(s,t)$, such that $\|\mathbf{x}-\mathbf{x}_{0}\|<q^{-\frac{s+t}{2}}$. By Taylor’s expansion (Lemma 2.2) of $\mathbf{f}$ around $\mathbf{x}_{0}$, the boundedness of second-order partial difference quotients of $f_{i}$ (i.e., (IV)), the ultrametric inequality (4.10), and (4.11), we obtain
$|Q_{0}+\mathbf{P}\cdot\mathbf{f}(\mathbf{x})|\leq\max\left\{q^{-t},q^{-\frac{s+t}{2}}q^{\frac{s-t}{2}},Mq^{s}\,\left(q^{-(\frac{s+t}{2})}\right)^{2}\right\}<\max\{1,M\}q^{-t}.$ (4.12)
Also, using the inequality (4.11) and Taylor’s expansion (Lemma 2.2) of $\partial_{i}f_{j}$ around the point $\mathbf{x}_{0}$, we obtain
$\left|\mathbf{P}_{i}^{(d)}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x})\mathbf{P}_{j}^{(m)}\right|$ $\leq$ $\max\left\{\left|\mathbf{P}_{i}^{(d)}+\sum_{j=1}^{m}\partial_{i}f_{j}(\mathbf{x}_{0})\mathbf{P}_{j}^{(m)}\right|,\max_{j=1,\dots,m}\left|\partial_{i}f_{j}(\mathbf{x})-\partial_{i}f_{j}(\mathbf{x}_{0})\right|\cdot\left|\mathbf{P}_{j}^{(m)}\right|\right\}$ (4.13)
$\leq$ $\max\left\{q^{\frac{s-t}{2}},Mq^{-\frac{s+t}{2}}q^{s}\right\}$
$\leq$ $\max\{1,M\}q^{\frac{s-t}{2}}.$
Finally, combining (4.12), (4.13), (4.10), and (4.11), there exists a constant $\widetilde{C}\in\mathbb{N}$ that depends only on $M$, $d$, and $n$, such that
$\begin{split}\mathfrak{M}(s,t)\subseteq\mathfrak{G}_{\mathbf{f}}\left(\widetilde{C}t,\widetilde{C}\frac{s-t}{2},\widetilde{C}s,\dots,\widetilde{C}s\right).\end{split}$
Lemma 4.7.
Let $U\subseteq\mathcal{K}^{d}$ be an open set and $\mathbf{f}:U\rightarrow\mathcal{K}^{n}$ be of the form $\mathbf{f}(\mathbf{x})=(\mathbf{x},\bm{f}(\mathbf{x}))$ such that it satisfies the hypothesis of Proposition 1. Then there exists $C,\alpha>0$ such that
$\mathcal{L}_{d}(\mathfrak{M}(s,t)\cap B)\leq C\max\left\{q^{-t},q^{\frac{(2n-1)s-3t}{2(n+1)}}\right\}^{\alpha}\mathcal{L}_{d}(B),$
whenever $t$ is large enough.
Proof.
Note that in the context of Theorem 4.5, up to a constant $\widetilde{C}$, we have
$r=t,s^{\prime}=\frac{s-t}{2},t_{1}=\cdots=t_{n}=s.$
Hence,
$\begin{split}s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}=(n-1)s-t+\frac{s-t}{2}=\frac{2n-1}{2}s-\frac{3}{2}t<0,\end{split}$
whenever $t$ is large enough. By Theorem 4.5 and Lemma 4.6, there exist $C>0$ and $\alpha>0$ such that
$\begin{split}\mathcal{L}_{d}\left(\mathfrak{M}(s,t)\cap B\right)\leq&C\max\left\{q^{-t},q^{\frac{(2n-1)s-3t}{2(n+1)}}\right\}^{\alpha}\mathcal{L}_{d}(B).\end{split}$
By combining Theorem 2.3, Lemma 4.3, and Lemma 4.7 we obtain Proposition 1.
5. Proof of Khintchine’s theorem: converegence case
Before proceeding to the proof of Theorem 1.2, we prove the following fact, which will be useful later in the proof of Theorem 1.2.
Lemma 5.1.
Let $\psi:(0,\infty)\to(0,1)$ be a non-increasing function. Then
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}<\infty\quad\Longleftrightarrow\quad\sum_{t=0}^{\infty}q^{t}\psi(q^{t})^{n}<\infty.$
Proof.
We know that every nonzero polynomial $Q\in\mathbb{F}_{q}[T]$ can be uniquely written as
$Q=a\cdot Q_{0},$
where $a\in\mathbb{F}_{q}^{\times}$ is a nonzero scalar and $Q_{0}$ is a monic polynomial. Furthermore, there are exactly $q^{t}$ monic polynomials of degree $t$, and each such $Q$ satisfies $|Q|=q^{t}$. Therefore,
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}=(q-1)\sum_{\begin{subarray}{c}Q\in\mathbb{F}_{q}[T]\setminus\{0\}\\ Q\text{monic}\end{subarray}}\psi(|Q|)^{n}=(q-1)\sum_{t=0}^{\infty}\sum_{\begin{subarray}{c}Q\text{ monic}\\ \deg Q=t\end{subarray}}\psi(q^{t})^{n}=(q-1)\sum_{t=0}^{\infty}q^{t}\psi(q^{t})^{n}.$
We are now in a position to prove Theorem 1.2.
Proof of Theorem 1.2.
Without loss of generality, it suffices to prove that for any $\mathbf{x}_{0}\in U$ and any small ball $B_{0}$ around $\mathbf{x}_{0}$
$\mathcal{L}_{d}\left(\left\{\mathbf{x}\in B_{0}:\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)\right\}\right)=0,$
whenever $\psi$ is monotonic and
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}<\infty.$
We divide the estimate into two parts: the special part, i.e., $\mathfrak{M}(-\log_{q}\psi(q^{t}),t)$, and the generic part, which comprises the set $B_{0}\setminus\mathfrak{M}(-\log_{q}\psi(q^{t}),t)$. Now we make the following observation. If $\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)$, then there exist infinitely many $t\in\mathbb{N}$ such that
$\left\|\mathbf{f}(\mathbf{x})-\frac{{\mathbf{P}}+\bm{\Theta}}{Q}\right\|\leq\frac{\psi(q^{t})}{q^{t}}\quad\text{with}\,\,(\mathbf{P},Q)\in\mathcal{R}^{n+1}\,\,\text{and}\,\,|Q|=q^{t}.$
Hence, for every $T\geq 1$, we have
$\begin{split}\left\{\mathbf{x}\in B_{0}:\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)\right\}\subseteq\bigcup_{t\geq T}\underbrace{\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})}_{A_{t}^{\bm{\Theta}}}\bigcup\\ \bigcup_{t\geq T}\underbrace{\bigcup_{\tiny(\mathbf{P},Q)\in\mathcal{R}_{\bm{\Theta}}(B_{0}\setminus\mathfrak{M}(-\log_{q}\psi(q^{t}),t))}\left\{\mathbf{x}\in B_{0}:\left\|\mathbf{x}-\frac{\pi(\mathbf{P}+\bm{\Theta})}{Q}\right\|\leq\frac{\psi(q^{t})}{q^{t}}\right\}}_{B_{t}^{\bm{\Theta}}},\end{split}$ (5.1)
where $\pi:\mathcal{K}^{n}\to\mathcal{K}^{d},$ is the projection into the first $d$ coordinates. To bound the measure of $B_{t}^{\bm{\Theta}}$, we invoke Proposition 2 to obtain
$\begin{split}\mathcal{L}_{d}(B_{t}^{\bm{\Theta}})\leq q^{t(d+1)}\psi(q^{t})^{m}\frac{\psi(q^{t})^{d}}{q^{td}}=q^{t}\psi(q^{t})^{n}.\end{split}$ (5.2)
By Lemma 5.1, the series $\sum_{t\geq T}\psi(q^{t})^{n}q^{t}$ converges. Therefore, by (5.2), $\lim_{T\rightarrow\infty}\mathcal{L}_{d}\left(\bigcup_{t\geq T}B_{t}^{\Theta}\right)=0$. For the other part, that is,
$A_{t}^{\bm{\Theta}}:=\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0},$
we apply Lemma 4.7 to estimate the measure of the set. Since $\lim_{|Q|\rightarrow\infty}\psi(|Q|)=0$, we may assume that $\psi(q^{t})\geq q^{-\frac{t}{2n-1}}$ for all $t$ sufficiently large. Otherwise, one can replace $\psi(q^{t})$ by $\max\{\psi(q^{t}),q^{-\frac{t}{2(n-1)}}\}$. Thus, $\log_{q}\psi(q^{t})\leq\frac{t}{2n-1}$ for every $t$ large enough. By Lemma 4.7, there exists $\alpha>0$ such that either
$\begin{split}\mathcal{L}_{d}\left(\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})\right)&\leq Cq^{\alpha\left(-\frac{2n-1}{2(n+1)}\log_{q}\psi(q^{t})-\frac{3}{2(n+1)}t\right)}\mathcal{L}_{d}(B_{0})\\ &\leq Cq^{-\frac{\alpha}{(n+1)}t}\mathcal{L}_{d}(B_{0}),\end{split}$ (5.3)
or,
$\mathcal{L}_{d}\left(\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})\right)\leq Cq^{-\alpha t}\mathcal{L}_{d}(B_{0}).$ (5.4)
Thus, by (5.2), (5.3) and (5.4), we have
$\sum_{t\geq T}\mathcal{L}_{d}(B_{t}^{\bm{\Theta}})+\sum_{t\geq T}\mathcal{L}_{d}(\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})<\infty,$
so that by the Borel-Cantelli lemma, $\mathcal{L}_{d}(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi)))=0$. ∎
5.1. Proof of Hausdorff Dimension
Proof of Theorem 1.3.
Take $s=-\log_{q}\psi(q^{t})>0$. By Theorem Proposition 1 and 2.3, one can cover $\mathfrak{M}(s,t)$ with disjoint balls of radius $q^{-\frac{s+t}{2}}$, such that the number of balls is $\ll q^{\frac{d}{2}(s+t)}\mathcal{L}_{d}(\mathfrak{M}(s,t))$. Thus by Proposition 1, there exists $T_{0}\geq 1$ such that for every $t\geq T_{0}$, we have
$\mathcal{H}^{\sigma}(A_{t}^{\bm{\Theta}})\ll q^{-\frac{s+t}{2}(\sigma-d)}\mathcal{L}_{d}(\mathfrak{M}(s,t))\ll q^{-\frac{s+t}{2}(\sigma-d)}.$ (5.5)
Hence, by plugging in $s=-\log_{q}\psi(q^{t})$,
[…]
For the set $B_{t}^{\bm{\Theta}}$, we employ the counting estimate (4.2) along with a covering by balls of radius $r=q^{-t}\psi(q^{t})$. Hence there exists $T_{1}\geq 1$ such that
$\sum_{t\geq T_{1}}\mathcal{H}^{\sigma}(B_{t}^{\bm{\Theta}})$ $\ll\sum_{t\geq T_{1}}q^{(d+1)t}q^{-ms}r^{\sigma}$
$=\sum_{t\geq T_{1}}{\psi(q^{t})}^{\sigma+m}q^{(d+1)t-\sigma t}$
[…]
Therefore, combining both for $t\geq T:=\max\{T_{0},T_{1}\}$, we obtain
$\sum_{t\geq T}\mathcal{H}^{\sigma}(A_{t}^{\bm{\Theta}})+\sum_{t\geq T}\mathcal{H}^{\sigma}(B_{t}^{\bm{\Theta}})<\infty.$ (5.7)
As a consequence, $\mathcal{H}^{\sigma}(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi)))=0$. ∎
6. Appendix - A Result in Geometry of Numbers for Discrete Subgroups
We consider an extension of the function field $\mathcal{K}$ obtained by adjoining the power $T^{\frac{1}{\ell}}$. As mentioned earlier, we denote this extension by $\mathcal{K}_{\ell}$. Similarly, denote the lattice $\mathbb{F}_{q}[T^{1/\ell}]\subseteq\mathcal{K}_{\ell}$ by $\widetilde{\mathcal{R}}:=\mathbb{F}_{q}[T^{\frac{1}{\ell}}]$. We now prove the following proposition, which establishes a relationship between the first minimum of the discrete subgroup $gu\mathcal{R}^{m+n}\subset\mathcal{K}_{\ell}^{m+n}$ and that of the lattice $gu\widetilde{\mathcal{R}}^{m+n}\subset\mathcal{K}_{\ell}^{m+n}$, where $g$ is some diagonal matrix in $\operatorname{SL}(m+n,\mathcal{K}_{\ell})$ and $u$ is some unipotent matrix in $\operatorname{SL}(m+n,\mathcal{K})$.
Proposition 3.
Let $g\in\operatorname{SL}\left(m+n,\mathcal{K}_{\ell}\right)$ be a diagonal matrix, and let $\bm{\alpha}\in M_{m\times n}(\mathcal{K})$ and
$u=\begin{pmatrix}\mathrm{I}_{m}&\bm{\alpha}\\ 0&I_{n}\end{pmatrix}\in\operatorname{SL}\left(m+n,\mathcal{K}\right).$
Then, $\lambda_{1}(gu\widetilde{\mathcal{R}}^{m+n})=\lambda_{1}(gu\mathcal{R}^{m+n})$.
Proof.
Let $(\mathbf{P},\mathbf{Q})=(P_{1},\dots,P_{m},Q_{1},\dots,Q_{n})^{T}\in\widetilde{\mathcal{R}}^{m+n}$ be a vector. Explicitly write $P_{i}=\sum_{k=0}^{l-1}T^{\frac{k}{\ell}}P_{i,k}$, where $P_{i,k}\in\mathcal{R}$, and $Q_{j}=\sum_{k=0}^{\ell-1}T^{\frac{k}{\ell}}Q_{j,k}$, where $Q_{j,k}\in\mathcal{R}$. Since every $\alpha\in\mathcal{K}_{\ell}$ can be written as $\alpha=T^{\deg(\alpha)}\cdot\frac{\alpha}{T^{\deg(\alpha)}}$, where $\deg(\alpha)\in\frac{1}{\ell}\mathbb{Z}$ and $\left|\frac{\alpha}{T^{\deg(\alpha)}}\right|=1$, there exist $r_{1},\dots,r_{m+n}\in\mathbb{Z}$ and $\mathbf{o}=\operatorname{diag}\{o_{1},\dots,o_{m+n}\}$, with $|o_{j}|=1$ for every $j=1,\dots,m+n$, such that
$g=\begin{pmatrix}T^{\frac{r_{1}}{\ell}}&&&\\ &\ddots&&\\ &&T^{\frac{r_{m+n}}{\ell}}&\end{pmatrix}\mathbf{o}.$
Since $\mathbf{o}$ does not affect the norm, we may assume that $\mathbf{o}=\mathrm{Id}$. Therefore,
$gu\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}=\begin{pmatrix}P_{1}+\sum_{j=1}^{n}\alpha_{1,j}Q_{j}\\ \vdots\\ P_{m}+\sum_{j=1}^{n}\alpha_{m,j}Q_{j}\\ Q_{1}\\ \vdots\\ Q_{n}\end{pmatrix}=\sum_{k=0}^{\ell-1}T^{\frac{k}{\ell}}\begin{pmatrix}P_{1,k}+\sum_{j=1}^{n}\alpha_{1,j}Q_{j,k}\\ \vdots\\ P_{m,k}+\sum_{j=1}^{n}\alpha_{m,j}Q_{j,k}\\ Q_{1,k}\\ \vdots\\ Q_{n,k}\end{pmatrix}$. (6.1)
Note that the norm of the $k$-th summand on the right-hand side of (6.1) is in $q^{\mathbb{Z}+\frac{k}{\ell}}$. Hence, every summand of each entry of (6.1) has a different absolute value or is equal to zero. Hence,
$\left\|gu\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}\right\|=\max_{k=0,\dots,\ell-1}q^{\frac{k}{\ell}}\left\|\begin{pmatrix}P_{1,k}+\sum_{j=1}^{n}\alpha_{1,j}Q_{j,k}\\ \vdots\\ P_{m,k}+\sum_{j=1}^{n}\alpha_{m,j}Q_{j,\ell}\\ Q_{1,k}\\ \vdots\\ Q_{n,k}\end{pmatrix}\right\|$. (6.2)
For $k=0,1,\dots,\ell-1$, let $\mathbf{P}^{(k)}=(P_{1,k},\dots,P_{m,k})^{T}$ and $\mathbf{Q}^{(k)}=(Q_{1,k},\dots,Q_{n,k})$. Let $k$ be such that
$\left\|gu\begin{pmatrix}\mathbf{P}^{(k)}\\ \mathbf{Q}^{(k)}\end{pmatrix}\right\|=\min\left\{\left\|gu\begin{pmatrix}\mathbf{P}^{(i)}\\ \mathbf{Q}^{(i)}\end{pmatrix}\right\|:gu\begin{pmatrix}\mathbf{P}^{(i)}\\ \mathbf{Q}^{(i)}\end{pmatrix}\neq 0\right\}.$
Then, by (6.2),
$\begin{split}\left\|gu\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}\right\|=\max_{i=0,\dots,\ell-1}q^{\frac{i}{\ell}}\left\|gu\begin{pmatrix}\mathbf{P}^{(i)}\\ \mathbf{Q}^{(i)}\end{pmatrix}\right\|\geq\left\|gu\begin{pmatrix}\mathbf{P}^{(k)}\\ \mathbf{Q}^{(k)}\end{pmatrix}\right\|\end{split}$ (6.3)
Furthermore $\begin{pmatrix}\mathbf{P}^{(k)}\\ \mathbf{Q}^{(k)}\end{pmatrix}\in\mathcal{R}^{m+n}$. As a consequence, there exists a vector $\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}\in\mathcal{R}^{m+n}$ such that
$\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)=\left\|gu\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}\right\|$.
Therefore, $\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)\geq\lambda_{1}\left(gu\mathcal{R}^{m+n}\right)$. On the other hand, $gu\mathcal{R}^{m+n}\subseteq gu\widetilde{\mathcal{R}}^{m+n}$, and thus, $\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)=\lambda_{1}\left(gu\mathcal{R}^{m+n}\right).$
Remark 6.1.
By Proposition 3, the set introduced in equation (4.1) can equivalently be written as
$\mathfrak{M}_{0}(s,t)=\left\{\mathbf{x}\in U:\lambda_{1}\!\left(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x})\mathcal{R}^{n+1}\right)=\lambda_{1}\!\left(g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x})\widetilde{\mathcal{R}}^{n+1}\right)<q^{-\frac{(d+2)t-2ns}{2(n+1)}}\right\}.$
This formulation offers an alternative characterization, and, most importantly, note that $g_{s,t}^{\star}u_{1}^{\star}(\mathbf{x})\widetilde{\mathcal{R}}^{n+1}$ is a lattice in $\mathcal{K}_{\ell}^{n+1}$. Hence, one can use the Duality theorem over the extended field $\mathcal{K}_{\ell}$ for the aforementioned lattice and have an alternative approach to provide a counting estimate for the generic part of the manifold $\mathscr{M}$.
Proposition 3 gives rise to several questions about the comparison between successive minima of discrete subgroups in extension fields and lattices in extension fields. Hence, we conclude this section with the following questions.
Question 6.2.
Let $\ell\geq 2$ and $n\in\mathbb{N}$. Also let $g\in\operatorname{SL}(n,\mathcal{K}_{\ell})$.
  1. (1)
    When do we have $\lambda_{1}(g\mathcal{R}^{n})=\lambda_{1}(g\widetilde{\mathcal{R}}^{n})$?
  2. (2)
    Given any $2\leq i\leq n$, when do we have $\lambda_{i}(g\mathcal{R}^{n})=\lambda_{i}(g\widetilde{\mathcal{R}}^{n})$?
  3. (3)
    Given $2\leq i\leq n$, when do we have $\lambda_{j}(g\mathcal{R}^{n})=\lambda_{j}(g\widetilde{\mathcal{R}}^{n})$ for every $j=1,\dots,i$?